OpenStudy (anonymous):

solve 6x+2y=0, -x-2y=-15

OpenStudy (anonymous):

Look at your previous question that I solved...You should look at people answers, and figure out how they get there, so you can do your homework...cause you can't ask anyone on the test...LOL

OpenStudy (anonymous):

i agree hanh! people here feel as if this place is a " do my homework" place. but we can help if you need it too.

OpenStudy (anonymous):

for -x-2y=-15, rearrange it you get x=15-2y. substitute it into 6x+2y=0, you get 6(15-2y)+2y=0. solve for it, get the value of y. sub back to x=15-2y and get your x value. try to understand the process. this will help when u have similar questions

OpenStudy (anonymous):

First add 15 to both sides of the 2nd equation, which gives you: -x - 2y + 15 = 0 6x + 2y = 0 From the equations above, you can set the two equations equal each other since then both = 0 So now you get: -x - 2y + 15 = 6x - 2y From there, add x and 2y to both sides, which then gives you 7x = 15 so therefore, x = 15/7 Now look at that and try to do similar problems so that you can past your class! >:)

OpenStudy (anonymous):

When you have x, substitute it back in to the two equations, and you should get the right answers for y.